# CALE : Concept-Aligned Embeddings for Both Within-Lemma and Inter-Lemma Sense Differentiation

### v1

@@ Abstract
Bastien Liétard1  and Gabriel Loiseau1,2, 1University of Lille, Inria, CNRS, Centrale Lille, UMR 9189 - CRIStAL, F-59000 Lille, France 2Hornetsecurity, Hem, France, first_name.last_name@inria.fr Authors contributed equally.
@@ 1 Introduction
Research in computational lexical semantics has relied on contextualized embeddings to study word meaning in context (Neidlein et al., 2020; Chronis and Erk, 2020; Apidianaki and Garí Soler, 2021; Yu and Xu, 2023; Li et al., 2024), with applications extending beyond traditional tasks, such as in debate modeling (Garí Soler et al., 2023) and political discourse analysis (Boholm et al., 2024). However, most approaches use pre-trained models like BERT (Devlin et al., 2019) or XLM-R (Conneau et al., 2020) without fine-tuning specifically for lexical semantics, due to limited annotated data.
While several models are fine-tuned for the Word-in-Context (WiC) task (Pilehvar and Camacho-Collados, 2019; Liu et al., 2021; Cassotti et al., 2023; Mosolova et al., 2024), WiC in its traditional definition only captures within-word meaning variation (i.e. the word’s senses) and ignores inter-word semantic relations. Consequently, models trained on WiC may struggle to capture the broader structure of lexical meaning.
To reach a higher perspective, the focus must shift from word senses to semantic concepts. Throughout this paper, we distinguish senses and concepts as follows: a word sense is conventionalized way to use a particular lemma in relevant contexts, it characterizes the use of a specific word for an intended meaning; concepts are mental representations of categories of objects, events, acts and ideas, and we use words to refer to them (Murphy, 2002). In context, the meaning of a word is the concept it refers to, and a word sense is a pointer to a concept. Our concepts are equivalent to Wikipedia entries, or WordNet synsets (Miller, 1995) as made in Liétard et al. (2024).
In this paper, we are interested in the use of contextualized models to obtain multi-purposes (as opposed to “task-specific”) vector representation of word meaning. Word occurrences that instantiate the same concept in respective contexts should have similar representations, while occurrences of words that refer to different concepts should have dissimilar embeddings. A model with this desired property would be useful in studies of the mapping between word forms and concepts, like Haspelmath (2023) for instance, or in any computational lexical semantic applications mentioned above.
@@
First, we introduce Concept Differentiation, a task that generalizes WiC to compare contextualized meanings of both same-lemma and cross-lemma word pairs. Given two occurrences, the task is to determine whether they instantiate the same semantic concept. Like WiC is to Word Sense Induction, Concept Differentiation is a binary classification form of Concept or Frame Induction (Liétard et al., 2024; QasemiZadeh et al., 2019).
@@
Then, we develop CALE (Concept-Aligned Embeddings), a family of token-level models fine-tuned for Concept Differentiation using SPCD. We compare them to their base pre-trained counterparts and to XL-LEXEME (Cassotti et al., 2023), a strong multilingual model trained on WiC-style data.
@@ 2.1 Lexical Semantic Tasks
A variety of NLP tasks have explored the mapping between words and meanings, especially regarding polysemy and synonymy (or similarity in meaning in general). Word Sense Disambiguation (WSD) maps word occurrences to predefined senses (Navigli, 2009; Bevilacqua et al., 2021), while its unsupervised counterpart, Word Sense Induction (WSI), clusters usages into latent sense groupings (Manandhar et al., 2010; Jurgens and Klapaftis, 2013). The Word-in-Context (WiC) task refines this by classifying whether two instances of a word share the same sense (Pilehvar and Camacho-Collados, 2019; Martelli et al., 2021). Lexical Semantic Change Detection (LSCD) extends WSI temporally, aiming to detect or rank meaning shifts across time periods (Schlechtweg et al., 2020; Zamora-Reina et al., 2022). While these tasks focus on polysemy and intra-word variation, they do not address inter-word semantic similarities. Tasks such as Lexical Substitution (McCarthy and Navigli, 2007; Zhou et al., 2019) and Lexical Similarity (Hill et al., 2015; Vulić et al., 2020; Huang et al., 2012; Armendariz et al., 2020) examine inter-word relations, centering on synonymy and contextual interchangeability. Substitution tasks often involve ranking appropriate alternatives in context (Erk and Padó, 2008; Kremer et al., 2014).
More recently, Concept Induction has been proposed to unify word-level and cross-word perspectives by clustering usages of multiple target words into semantic concepts (Liétard et al., 2024). Similarly, Frame Induction clusters word usages (typically verbs) into broader event frames with their associated arguments (QasemiZadeh et al., 2019; Yamada et al., 2021; Mosolova et al., 2024). These tasks aim to capture higher-order abstractions that go beyond individual word senses or pairwise similarity.
@@ 2.2 Embeddings for Semantic Representations
Contextualized word embeddings from models like ELMo (Peters et al., 2018) and BERT (Devlin et al., 2019) generate dynamic representations based on usage in context, enabling better semantic modeling than earlier static or sparse methods. However, pooling token-level embeddings into meaningful sentence-level vectors remains challenging due to issues like anisotropy in the embedding space (Reimers and Gurevych, 2019). To improve semantic similarity and retrieval, Sentence-BERT (SBERT) (Reimers and Gurevych, 2019) introduced a Siamese training architecture, followed by advances using contrastive learning (Gao et al., 2021) and multi-task objectives (Zhang et al., 2022) to enhance cross-task generalization. Such models have been adapted for lexical semantics. Mosolova et al. (2024) fine-tune language models using Wiktionary supervision to improve contextualized token embeddings for tasks like Word-in-Context. Similarly, the XL-LEXEME model (Cassotti et al., 2023), based on XLM-RoBERTa, extends SBERT for multilingual word meaning tasks, showing robustness on LSCD benchmarks and supporting cross-lingual, context-sensitive embedding spaces.
Other efforts fine-tune BERT for word-level classification tasks (Garí Soler and Apidianaki, 2020), though these cross-encoder models often lack flexibility for single-occurrence representation which represents a limitation for tasks requiring generalized concept modeling. Embedding-based models have also proven effective in capturing non-literal or stylistic nuances, as shown by applications in authorship verification (Rivera-Soto et al., 2021). Additionally, SenseBERT (Levine et al., 2020) enhances contextual embeddings by incorporating sense information from WordNet during pretraining, improving performance on sense-sensitive tasks and highlighting the benefit of explicitly modeling lexical semantics in transformer-based representations. Collectively, these developments demonstrate the adaptability of sentence-transformer architectures for fine-grained semantic representations.
@@ 3 Concept Differentiation
Our primary goal is to build a model able to accurately represent semantic information in order to distinguish concepts in context, from which we can obtain concept-aligned embeddings. To train such model, we rely on a supervised task that captures both polysemy and synonymy and propose Concept Differentiation, a supervised alternative to Concept Induction defined in Liétard et al. (2024).
@@ 4 Concept-Aligned Embeddings
We use our resulting pair classification dataset to train a representation model. The goal is to learn a function $f$ mapping an occurrence $o_{i}$ to a vector representation $e_{i}$ in a vector space in which the embeddings of two occurrences of the same concept have higher cosine similarity than embeddings of two occurrences of a different concept. To fit $f$, our architecture consists of a pre-trained contextualized language model followed by a pooling layer to obtain fixed-dimensional representations. We fine-tune this entire model using SentenceBERT-style siamese training and contrastive learning, which proceeds as follows.
@@
$\frac{1}{2}\left[y\cdot\mathrm{sim}_{ij}^{2}+(1-y)\cdot\max(0,m-\mathrm{sim}_{ij})^{2}\right]$
@@
using special tags <t></t> to delimit the target word in the sentence. During training, the model will learn that the contrast must be made based on the meaning of words between the delimiters. In doing so, we ensure the embedding corresponding to the full sequence is the representation of $o_{w,i}$ and distinct from the embedding of another target word occurring in the same sentence, without needing to account for potential subwords that may result from the model’s tokenizer. This method, also used by Cassotti et al. (2023), makes the model easy to use in practice at inference time, as we only have to add/move the delimiters to shift the model’s focus from one word to another in the sentence.
We refer to this fine-tuning approach and the resulting models as CALE (Concept Aligned Embeddings) and adopt this terminology throughout the paper. Our models are made publicly available through HuggingFace222[Hyperlink to HF-models]. We limited the scope of our study to single-word targets, but our framework could be applied to capture the meaning of Multi-Word Expressions in future work.
@@ 5 Model Evaluation
XLM-R (Conneau et al., 2020) is based on the RoBERTa Large architecture and pretrained on cross-lingual data. We consider three sets of layers to extract embeddings from : layers 1 to 4 (dubbed first), layers 14 to 17 (dubbed mid [^3] Chronis and Erk (2020) showed that late-intermediate layers were best for tasks related to semantic similarity) and the last four layers (21 to 24, dubbed last). Embeddings of word occurrences are obtained by averaging first over selected layers and then over subword tokens selection. XL-LEXEME (Cassotti et al., 2023) is a XLM-R model fine-tuned for multilingual WiC. It is a SentenceBERT model, meaning that a single embedding is produced per occurrence (using a final pooling layer). Compared to our models, XL-LEXEME’s fine-tuning was rather data-hungry, as three Word-in-Context datasets were used. XL-LEXEME established a new State-of-the-Art in some semantic tasks and is therefore our main reference point. Regrettably, models from Garí Soler and Apidianaki (2020) and Mosolova et al. (2024) are not publicly available, preventing us to include them in our experiments without re-implementing the whole dataset curation process and fine-tuning method. Our CALE models generate embeddings in a similar fashion than XL-LEXEME, but they are fine-tuned on Concept Differentiation, as described earlier. We experiment with several base model: XLM-R (to compare with XL-LEXEME) XL-LEXEME (to observe the impact of double fine-tuning and of XL-LEXEME’s fine-tuning multilinguality) and a monolingual English ModernBERT (Warner et al., 2024), a more-recent State-of-The-Art actualization of the BERT model, evaluated only in English datasets in order to discuss the potential difference between language specialization and pre-training multilinguality.
@@ 5.1 Concept Differentiation Evaluation
In Table 1 we present the results for the Concept Differentiation task on the test split of the SPCD datasets. For each model, the classification process for a given pair of occurrences is the following: we compute the embeddings for each occurrence using the candidate model, then compare the cosine distance between them to a threshold to label the pair: $1$ (same concept) if the cosine distance below the threshold, 0 (different concepts) otherwise. The threshold is chosen to maximize accuracy on the train and validation splits. We also provide a baseline system, where we label a pair as $1$ if and only if the two occurrences are from the same lemma. We refer to this baseline as 1L1C for “1 Lemma 1 Concept”. We use Balanced Accuracy (the average of the Recalls from both classes) as the classes are not exactly balanced. We also tried F1 scores and the observed tendencies are the same, thus we only report them in Appendix C, with also a more complete overview of accuracies per class.
@@
Our first observation is that the 1L1C baseline reaches high accuracy, providing a challenging baseline. This is in line with prior observations in the context of clustering (Liétard et al., 2024). This indicates that the mapping between words and concepts in SemCor may resemble a one-to-one mapping with some deviations, and as such a model for Concept Differentiation would benefit from not ignoring whether two occurrences are of the same lemma or not. We also observe that XLM-R performs best using late-intermediate layers than other layers and outperforms the baseline verifying findings of Chronis and Erk (2020) that these layers are best for tasks related to semantic similarity.
@@
To assess models’ ability to capture meaning variation within lemmas across contexts, we evaluate on the Lexical Semantic Change Detection (LSCD) task, following “Subtask 2“ of Schlechtweg et al. (2020). We use Diachronic Word Usage Graphs (DWUGs) datasets across six languages: English, German, Latin, Swedish, Spanish, and Italian (Schlechtweg et al., 2020; Zamora-Reina et al., 2022; Cassotti et al., 2024) [^4] https://www.ims.uni-stuttgart.de/en/research/resources/experiment-data/wugs/. Each dataset contains target words with occurrences sampled from two time periods ($T_{1}$, $T_{2}$), and the task is to assign a change score reflecting how much a word’s meaning has shifted. LSCD is a intra-lemma task: only usages of the same word are compared and semantic change is assessed by measuring differences in the word’s senses between time periods. Performance is measured by Spearman correlation between predicted and gold change scores. Note that Cassotti et al. (2023) also evaluated their model XL-LEXEME on these benchmarks and established new State-of-the-Art scores in English, German and Swedish.
We use the Average Pairwise Divergence (APD) measure from Kutuzov and Giulianelli (2020) to compute change scores, relying on cosine distances between embeddings from $T_{1}$ and $T_{2}$:
@@
with $d$ the cosine distance operator, $e_{w,i}^{k}$ the embedding of the $i$-th occurrence of word $w$ at time $T_{k}$, for a target word $w$ with $n$ occurrences at $T_{1}$ and $m$ occurrences at $T_{2}$. APD is generally accepted as standard for Lexical Semantic Change (see the survey study of Periti and Montanelli (2024)) and was used by Cassotti et al. (2023) to evaluate XL-LEXEME on the same task on DWUGs datasets. In Appendix D we tried PRT, another measure from Kutuzov and Giulianelli (2020).
@@
Subtask 1 : Contextual Changes
EN FI HR SL
# examples 340 24 112 111
XLM-R (first) .18 .32 .20 .33
XLM-R (mid) .69 .36 .58 .52
XLM-R (last) .63 .27 .51 .45
XL-LEXEME .70 .47 .67 .66
CALE (XLM-R) .74 .63 .72 .70
CALE (MBERT) .68 - - -
CALE (XL-LEX) .74 .60 .72 .67
Subtask 2 : Similarity Ratings
EN FI HR SL
# pairs 680 48 224 222
XLM-R (first) .27 -.04 .25 .25
XLM-R (mid) .61 .25 .53 .51
XLM-R (last) .59 .16 .52 .46
XL-LEXEME .63 .59 .64 .67
CALE (XLM-R) .67 .64 .68 .68
CALE (MBERT) .67 - - -
CALE (XL-LEX) .67 .57 .73 .68
@@
We now evaluate candidate models in the cross-lemma setting. To assess our models’ ability to represent conceptual similarity of two distinct words in a given context, we use the CoSimLex dataset (Armendariz et al., 2020) for In-Context Lexical Similarity. This resource proposes two subtasks in 4 different languages, English, Finnish, Croatian and Slovene. Each entry of CoSimLex is a pair of target words, and two different contexts in which both words appear. For each context is provided a measure of lexical similarity between the two words. Subtask 1, Contextual Changes, aims at predicting the change of similarity between the two words, from context 1 to context 2. This is measured with Pearson Correlation coefficient between the differences of gold similarities and the differences of predicted similarities. Subtask 2, Similarity Ratings, adopts a more traditional setting, where candidate systems are simply tasked to predict a measure of lexical similarity between the two words in each context, and measuring the Spearman Correlation between gold measures of similarity and predicted ones. Subtask 2 uses independently each context provided in Subtask 1 instead of pairing them, and as such covers twice as many examples. To predict lexical similarity between the target lemmas in a given context for both subtasks, we compute the embeddings corresponding to each target and compute the cosine similarity between them. Results are displayed in Table 3.
CALE models reach the highest correlations values in all languages, ranging between $0.63$ to $0.73$ in subtask 1 and $0.57$ to $0.68$ in subtask 2, and consistently improve over XLM-R. Surprisingly, XL-LEXEME also performs very well and is most often not deemed significantly different from CALE results, despite being trained only on a same-lemma task and not on inter-words relations. CALE (XL-LEX) does not perform significantly better than the more simple CALE (XLM-R), indicating that CALE fine-tuning is usually enough to obtain accurate in-context similarities. Again, the good performances reached by CALE fine-tuning highlight the cross-lingual transferability of XLM-R models.

### v2

@@ Abstract
Bastien Liétard1  and Gabriel Loiseau1,2, 1University of Lille, Inria, CNRS, Centrale Lille, UMR 9189 - CRIStAL, F-59000 Lille, France 2Hornetsecurity, Hem, France first_name.last_name@inria.fr Authors contributed equally.
@@ 1 Introduction
Research in computational lexical semantics has relied on contextualized embeddings to study word meaning in context (neidlein-etal-2020-analysis; chronis-erk-2020-bishop; apidianaki-gari-soler-2021-dolphins; yu-xu-2023-word; li-etal-2024-investigating), with applications extending beyond traditional tasks, such as in debate modeling (gari-soler-etal-2023-measuring) and political discourse analysis (boholm-etal-2024-political). However, most approaches use pre-trained models like BERT (devlin-etal-2019-bert) or XLM-R (conneau-etal-2020-unsupervised) without fine-tuning specifically for lexical semantics, due to limited annotated data.
While several models are fine-tuned for the Word-in-Context (WiC) task (pilehvar-camacho-collados-2019-wic; liu-etal-2021-mirrorwic; cassotti-etal-2023-xl; mosolova-etal-2024-injecting), WiC in its traditional definition only captures within-word meaning variation (i.e. the word’s senses) and ignores inter-word semantic relations. Consequently, models trained on WiC may struggle to capture the broader structure of lexical meaning.
To reach a higher perspective, the focus must shift from word senses to semantic concepts. Throughout this paper, we distinguish senses and concepts as follows: a word sense is a conventionalized way to use a particular lemma in relevant contexts, it characterizes the use of a specific word for an intended meaning; concepts are mental representations of categories of objects, events, acts and ideas, and we use words to refer to them (Murphy, 2002). In context, the meaning of a word is the concept it refers to, and a word sense is a pointer to a concept. Our concepts are equivalent to Wikipedia entries, or WordNet synsets (Miller, 1995) as made in lietard-etal-2024-word.
In this paper, we are interested in the use of contextualized models to obtain multi-purpose (as opposed to “task-specific”) vector representation of word meaning. Word occurrences that instantiate the same concept in respective contexts should have similar representations, while occurrences of words that refer to different concepts should have dissimilar embeddings. A model with this desired property would be useful in studies of the mapping between word forms and concepts, like Haspelmath (2023) for instance, or in any computational lexical semantic applications mentioned above.
@@
First, we introduce Concept Differentiation, a task that generalizes WiC to compare contextualized meanings of both same-lemma and cross-lemma word pairs. Given two occurrences, the task is to determine whether they instantiate the same semantic concept. Like WiC is to Word Sense Induction, Concept Differentiation is a binary classification form of Concept or Frame Induction (lietard-etal-2024-word; qasemizadeh-etal-2019-semeval).
@@
Then, we develop CALE (Concept-Aligned Embeddings), a family of token-level models fine-tuned for Concept Differentiation using SPCD. We compare them to their base pre-trained counterparts and to XL-LEXEME (cassotti-etal-2023-xl), a strong multilingual model trained on WiC-style data.
@@ 2.1 Lexical Semantic Tasks
A variety of NLP tasks have explored the mapping between words and meanings, especially regarding polysemy and synonymy (or similarity in meaning in general). Word Sense Disambiguation (WSD) maps word occurrences to predefined senses (Navigli, 2009; Bevilacqua et al., 2021), while its unsupervised counterpart, Word Sense Induction (WSI), clusters usages into latent sense groupings (manandhar-etal-2010-semeval; jurgens-klapaftis-2013-semeval). The Word-in-Context (WiC) task refines this by classifying whether two instances of a word share the same sense (pilehvar-camacho-collados-2019-wic; martelli-etal-2021-semeval). Lexical Semantic Change Detection (LSCD) extends WSI temporally, aiming to detect or rank meaning shifts across time periods (schlechtweg-etal-2020-semeval; d-zamora-reina-etal-2022-black). While these tasks focus on polysemy and intra-word variation, they do not address inter-word semantic similarities. Tasks such as Lexical Substitution (mccarthy-navigli-2007-semeval; zhou-etal-2019-bert) and Lexical Similarity (hill-etal-2015-simlex; vulic-etal-2020-multi; huang-etal-2012-improving; armendariz-etal-2020-cosimlex) examine inter-word relations, centering on synonymy and contextual interchangeability. Substitution tasks often involve ranking appropriate alternatives in context (erk-pado-2008-structured; kremer-etal-2014-substitutes).
More recently, Concept Induction has been proposed to unify word-level and cross-word perspectives by clustering usages of multiple target words into semantic concepts (lietard-etal-2024-word). Similarly, Frame Induction clusters word usages (typically verbs) into broader event frames with their associated arguments (qasemizadeh-etal-2019-semeval; yamada-etal-2021-semantic; mosolova-etal-2024-injecting). These tasks aim to capture higher-order abstractions that go beyond individual word senses or pairwise similarity.
@@ 2.2 Embeddings for Semantic Representations
Contextualized word embeddings from models like ELMo (peters-etal-2018-deep) and BERT (devlin-etal-2019-bert) generate dynamic representations based on usage in context, enabling better semantic modeling than earlier static or sparse methods. However, pooling token-level embeddings into meaningful sentence-level vectors remains challenging due to issues like anisotropy in the embedding space111i.e., the tendency for embeddings to occupy a narrow cone, leading to high cosine similarity for even unrelated items (Gao et al., 2019). (reimers-gurevych-2019-sentence). To improve semantic similarity and retrieval, Sentence-BERT (SBERT) (reimers-gurevych-2019-sentence) introduced a Siamese training architecture, followed by advances using contrastive learning (gao-etal-2021-simcse) and multi-task objectives (zhang-etal-2022-task) to enhance cross-task generalization. Such models have been adapted for lexical semantics. mosolova-etal-2024-injecting fine-tune language models using Wiktionary supervision to improve contextualized token embeddings for tasks like Word-in-Context. Similarly, the XL-LEXEME model (cassotti-etal-2023-xl), based on XLM-RoBERTa (conneau-etal-2020-unsupervised), extends SBERT for multilingual word meaning tasks, showing robustness on LSCD benchmarks and supporting cross-lingual, context-sensitive embedding spaces.
Other efforts fine-tune BERT for word-level classification tasks (gari-soler-apidianaki-2020-multisem), though these cross-encoder models often lack flexibility for single-occurrence representation which represents a limitation for tasks requiring generalized concept modeling. Embedding-based models have also proven effective in capturing non-literal or stylistic nuances, as shown by applications in authorship verification (rivera-soto-etal-2021-learning). Additionally, SenseBERT (levine-etal-2020-sensebert) enhances contextual embeddings by incorporating sense information from WordNet during pretraining, improving performance on sense-sensitive tasks and highlighting the benefit of explicitly modeling lexical semantics in transformer-based representations. Collectively, these developments demonstrate the adaptability of sentence-transformer architectures for fine-grained semantic representations.
@@ 3 Concept Differentiation
Our primary goal is to build a model able to accurately represent semantic information in order to distinguish concepts in context, from which we can obtain concept-aligned embeddings. To train such model, we rely on a supervised task that captures both polysemy and synonymy and propose Concept Differentiation, a supervised alternative to Concept Induction defined in lietard-etal-2024-word.
@@ 4 Concept-Aligned Embeddings
We use our resulting pair classification dataset to train a representation model. The goal is to learn a function $f$ mapping an occurrence $o_{i}$ to a vector representation $e_{i}$ in a vector space in which the embeddings of two occurrences of the same concept have higher cosine similarity than embeddings of two occurrences of a different concept. To fit $f$, we adopt a Siamese representation learning architecture, inspired from SentenceBERT. Concretely, our model consists of a pre-trained contextualized language model followed by a pooling layer that produces a fixed-dimensional vector for an input sequence. During training, two identical copies of this encoder sharing the same parameters are used to process two occurrences as a training pair. The parameters of the encoder are updated jointly, ensuring that both occurrences are embedded in the same representation space.
@@
$\frac{1}{2}\left[y\cdot\operatorname{sim}_{ij}^{2}+(1-y)\cdot\max(0,m-\operatorname{sim}_{ij})^{2}\right]$
@@
using special tags <t></t> to delimit the target word in the sentence. During training, the model will learn that the contrast must be made based on the meaning of words between the delimiters. In doing so, we ensure the embedding corresponding to the full sequence is the representation of $o_{w,i}$ and distinct from the embedding of another target word occurring in the same sentence, without needing to account for potential subwords that may result from the model’s tokenizer. This method, also used by cassotti-etal-2023-xl, makes the model easy to use in practice at inference time, as we only have to add/move the delimiters to shift the model’s focus from one word to another in the sentence.
We refer to this fine-tuning approach and the resulting models as CALE (Concept Aligned Embeddings) and adopt this terminology throughout the paper. Our models are made publicly available through HuggingFace333[Hyperlink to HF-models]. We limited the scope of our study to single-word targets, but our framework could be applied to capture the meaning of Multi-Word Expressions in future work.
@@ 5 Model Evaluation
XLM-R (conneau-etal-2020-unsupervised) is based on the RoBERTa Large architecture and pretrained on cross-lingual data. XL-LEXEME (cassotti-etal-2023-xl) is a XLM-R model fine-tuned for multilingual WiC. Compared to our models, XL-LEXEME’s fine-tuning was rather data-hungry, as three Word-in-Context datasets were used. XL-LEXEME established a new State-of-the-Art in some semantic tasks and is therefore our main reference point. Regrettably, models from gari-soler-apidianaki-2020-multisem and mosolova-etal-2024-injecting are not publicly available, preventing us to include them in our experiments without re-implementing the whole dataset curation process and fine-tuning method. Our CALE models generate embeddings in a similar fashion than XL-LEXEME, but they are fine-tuned on Concept Differentiation, as described earlier. We experiment with several base model: XLM-R (to compare with XL-LEXEME) XL-LEXEME (to observe the impact of double fine-tuning and of XL-LEXEME’s fine-tuning multilingually) and a monolingual English ModernBERT (Warner et al., 2024), a more-recent State-of-The-Art actualization of the BERT model, evaluated only in English datasets in order to discuss the potential difference between language specialization and pre-training multilingually.
Embedding extraction.
For XLM-R, embeddings of word occurrences are obtained by averaging first over a specified range of layers, and then over subword tokens that compose the target word of the occurrence. XL-LEXEME and CALE models are SentenceBERT models, meaning that a single embedding is produced per occurrence using a final pooling layer that averages the last layer embeddings of all tokens in the sequence. In that case, the delimiters around the target word ensure that the representation is specific to that target.
Hyperparameters and metrics.
For XLM-R, we test three layer ranges. The mid layer set corresponds to intermediate-high layers (14 to 17), following the findings of chronis-erk-2020-bishop; the last set (layers 21 to 24) is in line with a large number of works in lexical semantic tasks using the average over the last four layers as input; and first (layers 1 to 4) is included for comprehensiveness. For all models and across all experiments, cosine distance is used as the standard metric of the representational space to compare two embeddings.
@@ 5.1 Concept Differentiation Evaluation
In Table 1 we present the results for the Concept Differentiation task on the test split of the SPCD datasets.
Threshold-based classifiers.
For each candidate model, the classification process for a given pair of occurrences is the following: we compute the embeddings for each occurrence using the candidate model, then compare the cosine distance between them to a threshold to label the pair: $1$ (same concept) if the cosine distance below the threshold, 0 (different concepts) otherwise. The threshold is chosen to maximize accuracy on the train and validation splits.
Baseline.
To provide a baseline system, we label a pair as $1$ if and only if the two occurrences are from the same lemma, and 0 otherwise. We refer to this baseline as 1L1C for “1 Lemma 1 Concept”.
Evaluation metrics.
We use Balanced Accuracy (the average of the Recalls from both classes) as the classes are not exactly balanced. We also tried F1 scores and the observed tendencies are the same, thus we only report them in Appendix C, with also a more complete overview of accuracies per class.
@@
Discussion.
Our first observation is that the 1L1C baseline reaches high accuracy, providing a challenging baseline. This is in line with prior observations in the context of clustering (lietard-etal-2024-word). This indicates that the mapping between words and concepts in SemCor may resemble a one-to-one mapping with some deviations, and as such a model for Concept Differentiation would benefit from not ignoring whether two occurrences are of the same lemma or not. We also observe that XLM-R performs best using late-intermediate layers than other layers and outperforms the baseline verifying findings of chronis-erk-2020-bishop that these layers are best for tasks related to semantic similarity.
@@ Discussion.
To assess models’ ability to capture meaning variation within lemmas across contexts, we evaluate on the Lexical Semantic Change Detection (LSCD) task, following “Subtask 2“ of schlechtweg-etal-2020-semeval. We use Diachronic Word Usage Graphs (DWUGs) datasets across six languages: English, German, Latin, Swedish, Spanish, and Italian (schlechtweg-etal-2020-semeval; d-zamora-reina-etal-2022-black; cassotti-etal-2024-dwugs) [^4] https://www.ims.uni-stuttgart.de/en/research/resources/experiment-data/wugs/. Each dataset contains target words with occurrences sampled from two time periods ($T_{1}$, $T_{2}$), and the task is to assign a change score reflecting how much a word’s meaning has shifted. LSCD is a intra-lemma task: only usages of the same word are compared and semantic change is assessed by measuring differences in the word’s senses between time periods. Performance is measured by Spearman correlation between predicted and gold change scores. Note that cassotti-etal-2023-xl also evaluated their model XL-LEXEME on these benchmarks and established new State-of-the-Art scores in English, German and Swedish.
We use the Average Pairwise Divergence (APD) measure from kutuzov-giulianelli-2020-uio to compute change scores, relying on cosine distances between embeddings from $T_{1}$ and $T_{2}$:
@@
with $d$ the cosine distance operator, $e_{w,i}^{k}$ the embedding of the $i$-th occurrence of word $w$ at time $T_{k}$, for a target word $w$ with $n$ occurrences at $T_{1}$ and $m$ occurrences at $T_{2}$. APD is generally accepted as standard for Lexical Semantic Change (see the survey study of Periti and Montanelli (2024)) and was used by cassotti-etal-2023-xl to evaluate XL-LEXEME on the same task on DWUGs datasets. In Appendix D we tried PRT, another measure from kutuzov-giulianelli-2020-uio.
@@
Context Similarity [][...] As J.T. put it, if Kaz can’t follow orders, he might as well learn to give them. 2.88 [][...] Caleb returns to the table to let Kirsten know that he will give his support to Julie at the board meeting. [][...] but for her marriage’s sake, she will put her support to Sandy. 7.83
Table 3: Example entry of CoSimLex context pairs with the mean similarity rating between target words (underlined).
Subtask 1 : Contextual Changes EN FI HR SL # examples 340 24 112 111 XLM-R (first) .18 .32 .20 .33 XLM-R (mid) .69 .36 .58 .52 XLM-R (last) .63 .27 .51 .45 XL-LEXEME .70 .47 .67 .66 CALE (XLM-R) .74 .63 .72 .70 CALE (MBERT) .68 - - - CALE (XL-LEX) .74 .60 .72 .67 Subtask 2 : Similarity Ratings EN FI HR SL # pairs 680 48 224 222 XLM-R (first) .27 -.04 .25 .25 XLM-R (mid) .61 .25 .53 .51 XLM-R (last) .59 .16 .52 .46 XL-LEXEME .63 .59 .64 .67 CALE (XLM-R) .67 .64 .68 .68 CALE (MBERT) .67 - - - CALE (XL-LEX) .67 .57 .73 .68
@@
We now evaluate candidate models in the cross-lemma setting.
Dataset.
To assess our models’ ability to represent conceptual similarity of two distinct words in a given context, we use the CoSimLex dataset (armendariz-etal-2020-cosimlex) for In-Context Lexical Similarity. This resource proposes two subtasks in 4 different languages, English, Finnish, Croatian and Slovene. Each entry of CoSimLex is a pair of target words, and two different contexts in which both words appear. For each context is provided a measure of lexical similarity between the two words. An example of a CoSimLex entry in English is provided in Table 3, where the respective meanings of put and give are closer in context 2 than in context 1.
Tasks.
Subtask 1, Contextual Changes, aims at predicting the change of similarity between the two words, from context 1 to context 2. This is measured with Pearson Correlation coefficient between the differences of gold similarities and the differences of predicted similarities. Subtask 2, Similarity Ratings, adopts a more traditional setting, where candidate systems are simply tasked to predict a measure of lexical similarity between the two words in each context, and measuring the Spearman Correlation between gold measures of similarity and predicted ones. Subtask 2 uses independently each context provided in Subtask 1 instead of pairing them, and as such covers twice as many examples.
Predicting lexical similarity in the same-context setting.
For both subtasks, given a context from CoSimLex where two target words appear, we compute the embeddings corresponding to each target and then compare them using cosine distance. For XLM-R, we take the embeddings corresponding to the two sets of subwords, that compose each target. For XL-LEXEME and CALE, we create two separate input sequences: the first input contains the full context with our delimiter tags placed around the first target word only; and the second with delimiters around the second target only. Feeding each sequence to the model, we obtain two embeddings, each representing the meaning of the specifically tagged target word.
Results and discussion.
Results are displayed in Table 4. CALE models reach the highest correlations values in all languages, ranging between $0.63$ to $0.73$ in subtask 1 and $0.57$ to $0.68$ in subtask 2, and consistently improve over XLM-R. Surprisingly, XL-LEXEME also performs very well and is most often not deemed significantly different from CALE results, despite being trained only on a same-lemma task and not on inter-words relations. CALE (XL-LEX) does not perform significantly better than the more simple CALE (XLM-R), indicating that CALE fine-tuning is usually enough to obtain accurate in-context similarities. Again, the good performances reached by CALE fine-tuning highlight the cross-lingual transferability of XLM-R models.